Make a the subject of the formula in s=n\2[2a+(n-1)d]

Answer by Boreal(15235)

Solve for a.
Isolate a
get rid of the fraction
2s=n (2a+(n-1)d)
(2s/n)-(n-1)d=2a
(s/n)-[(n-1)d]/2=a
--
check by using the arithmetic series (the formula is for the sum of one) 1,3,5,7,9,11,13,15
s=(8/2)(2+(7(2))
=4(16)
=64
-
now solve for the first term
(64/8)-(7*2)/2=8-7=1